The air show has 5 more biplanes and gliders. Together there are 23 flying machines. How many biplanes are there?
Answer by stanbon(75887)

The air show has 5 more biplanes then gliders. Together there are 23 flying machines. How many biplanes are there?
----
Let # of gliders be "x".
Then # of biplanes is "x+5".
-----
Equation:
x + x+5 = 23
2x = 18
x = 9 (# of gliders)
x+5 = 14 (3 of biplanes)
---------
Cheers,
Stan H.